OpenStudy (anonymous):

Find the zeros in ; x^3-6x^2-9x+54

OpenStudy (bahrom7893):

x^2(x-6) -9 (x-6) = (x-6)(x^2-9)=(x-6)(x-3)(x+3)=0

OpenStudy (bahrom7893):

Try that

OpenStudy (anonymous):

It works thank you